OpenStudy (anonymous):

Two factory plants are making TV panels. Yesterday, Plant A produced 3000 fewer panels than Plant B did. Three percent of the panels from Plant A and 5% of the panels from Plant B were defective. How many panels did Plant B produce, if the two plants together produced 1190 defective panels?

OpenStudy (anonymous):

Production from plant a = (b-3000) 3% of a plus 5% of b = 1190 so, (b-3000) x 3/100 + b x 5/100 = 1190 From this you can work out b

OpenStudy (anonymous):

so to get b i do 8% of 1190?

OpenStudy (anonymous):

it will be 0.03b-90+0.05b=1190 then when you combine like terms it turns into 0.08b-90=1190 then you can just solve for b by dividing 1190 by 0.08 so technically yes you are right haha i just had to write out the math real fast